Convert each number into scientific notation

  1. 4,000,000,000
  2. 0.000000027
  3. 67,760,000
  4. 90,500
  5. 540,900
  6. 0.00904
  7. 0.032
  8. 512,000,000
Answer #1

Scientific notation is a way of writing very big numbers or very small numbers in terms of powers of 10.

The number is written in two parts:

● The digits with the decimal point after the first digit

● followed by × 10 to a power that puts the decimal point where it should be

Note that value of the number doesn't changes at all only the representation of number gets changed.

For example:- 65000000 can be written as 6.5× 107 in scientific notation form.

Therefore,

4,000,000,000= 4×109

0.000000027= 2.7 × 10-8

67,760,000=6.776×107

90,500=9.05×104

540,900=5.409×105

0.00904=9.04×10-3

0.032=3.2×10-2

512,000,000=5.12×108

Hope you will like my answer:)